OpenStudy (anonymous):

On the first day of the year, Diego had $700 in his savings account and started spending $35 a week. His brother Juan had $450 and started saving $15 a week. After how many weeks will the brothers have the same amount? What will that amount be? Help !

OpenStudy (anonymous):

700 - 35 every week 450 + 15 every week does that help? you can solve it yourself, try!

OpenStudy (anonymous):

do i put them together

OpenStudy (anonymous):

x is the number of weeks Diego's amount after x weeks: 700-35x Juan's amount after x weeks: 450 + 15x These are equal so write equation 700-35x =450 +15x 700-450 = 35x +15x 250 = 50x x = 5 weeks 450 +15*5 = 450 +75 = $525

OpenStudy (anonymous):

thanks